1. Write the balanced NET ionic equation for the reaction when Li₂SO₄ and Ca(NO₃)₂ are mixed in aqueous solution. If no reaction occurs, write only NR. 2. Write the balanced NET ionic equation for the reaction when aluminum nitrate and potassium iodide are mixed in aqueous solution. If no reaction occurs, write only NR. 3. Write the balanced NET ionic equation for the reaction when Al(NO₃)₃ and Na₃PO₄ are mixed in aqueous solution. If no reaction occurs, write only NR.
1
Expert's answer
2019-02-19T04:07:48-0500

Answer:

1) Ca(NO3)2 + Li2SO4 = 2LiNO3 + CaSO4

Ca2+ + 2NO3- + 2Li+ + SO42- = 2Li+ + 2NO3- + CaSO4

Ca2+ + SO42- = CaSO4


2) Al(NO3)3 + 3KI = AlI3 + 3KNO3 - reversible process


3) Al(NO3)3 + Na3PO4 = AlPO4 + 3NaNO3

Al3+ + 3NO3- + 3Na+ + PO43- = AlPO4 + 3Na+ + 3NO3-

Al3+ + PO43- = AlPO4